# Analysis of the $q\bar{q}\to Z^{*}\to hA\to 4\tau$ process within the lepton-specific 2HDM at the LHC

### v2

@@ Abstract
We analyse light Higgs scalar and pseudoscalar associated hadro-production in the 2-Higgs Doublet Model (2HDM) Type-X (or lepton-specific) within the parameter space allowed by theoretical self-consistency requirements as well as the latest experimental constraints from the Large Hadron Collider (LHC), precision data and $B$ physics. Over the viable regions of such a scenario, the Standard Model-like Higgs boson discovered at the LHC in 2012 is the heavier CP-even state $H$. Furthermore, in the Type-X scenario, due to large $\tan\beta$, the lighter Higgs scalar $h$ and the pseudoscalar $A$ mainly decay into two $\tau$ leptons. Therefore, we concentrate on analysing the signal process $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}\to\ell\nu_{\ell}\ell\nu_{\ell}\tau_{h}\tau_{h}$ (where $\ell=e,\mu$ whereas $\tau_{h}$ represents the hadronic decay of the $\tau$) and explore the feasibility of conducting such a search at the LHC with a centre-of-mass energy of $\sqrt{s}~{}=~{}\text{14 TeV}$ and a luminosity of $L~{}=~{}\text{300}~{}\text{fb}^{-1}$. To suppress the huge SM background, we confine ourselves to consider the fraction of signal events with two same-sign $\tau$ leptons further decaying into same-sign leptons while the other two $\tau$ leptons decay hadronically. We find that a combination of kinematical selection and machine learning (ML) analysis will yields significant sensitivity to this process at the end of the LHC Run 3.
@@ I Introduction
The LHC has played an indispensable role in improving our understanding of the microcosmic subatomic structure, particularly with the discovery of a 125 GeV scalar particle by the ATLAS and CMS experimental teams in 2012 Aad:2012tfa ; Chatrchyan:2012ufa . This was the last particle predicted by the Standard Model (SM) and its properties measured in the above experiments are currently consistent with the SM prediction at the $5~{}\sigma$ level ATLAS:2022vkf ; CMS:2022dwd . Investigating the Higgs sector at the LHC, where the presence of additional (pseudo)scalars is possible, remains an intriguing task, particularly in new physics models with additional doublet fields. Therefore, to discover extra Higgs bosons at the LHC is one of the prime targets of Beyond the SM (BSM) signal searches.
@@
One of the simplest extensions of the SM is the 2HDM, which contains two complex Higgs doublets Lee:1973iz ; Deshpande:1977rw ; Branco:2011iw . After Electro-Weak Symmetry Breaking (EWSB), there are three Goldstone bosons, which are ‘eaten’ by the $W^{\pm}$ and ${Z}$ bosons while the remaining five degrees of freedom incarnate into physical Higgs bosons: two neutral CP-even scalars (${h}$ and ${H}$, with $M_{{h}}\textless M_{{H}}$), one CP-odd pseudoscalar (${A}$) and two charged Higgs states ${H}^{\pm}$ (with mixed CP status).
@@
In order to be consistent with the stringent experimental constraints from Flavour Changing Neutral Currents (FCNCs) at the tree level, typically, a $Z_{2}$ symmetry is introduced into the Yukawa sector so that each type of fermion only couples to one of the doublets of the 2HDM Glashow:1976nt . Based on the $Z_{2}$ charge assignments of the Higgs doublets, we can define four basic scenarios, known as (Yukawa) types. In the lepton-specific scenario (or Type-X), the additional Higgs doublet $\phi_{2}$ only couples to the lepton sector and not to the quark sector, where the light Higgs boson mainly decays to $\tau^{+}\tau^{-}$ within a wide range of parameter spaces Aoki:2009ha .
@@
Motivated by recent searches for $\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ by the CMS experiment CMS:2015twz ; CMS:2017dmg ; CMS:2019spf , we will study the process $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ in the 2HDM Type-X. In this scenario, the $h/A\to\tau^{+}\tau^{-}$ decay could be enhanced, resulting in a sizeable $4\tau$ final state. Therefore, it is worthwhile to examine whether this statement is robust enough after taking into account the effects of parton shower, hadronisation, heavy flavour decays and detector resolution.
@@
The recent results by the (muon) $g-2$ experiment conducted by the Fermilab National Accelerator Laboratory (FNAL), when combined with the old ones from the Brookhaven National Laboratory (BNL), yield a 4.2 $\sigma$ discrepancy between the experimental data and the theoretical prediction from the SM:
$\Delta a_{\mu}=a_{\mu}^{\exp}-a_{\mu}^{\rm SM}=(25.1\pm 5.9)\times 10^{10},$ (1)
@@
where $a_{\mu}=(g-2)_{\mu}/2$  Albahri2021 . One can explain this using 2HDM Type-X model, though, since the couplings of the new Higgs bosons to the SM fermions are proportional to $\text{tan}\beta$, thereby enhancing the corresponding contributions to $\Delta a_{\mu}$ Jueid2023 ; Iguro:2023tbk . Although a recent measurement has already shown possibilities of restricting the parameter spaces of the Type-X model in the large $\text{tan}\beta$ region CMS:CMS23007 , further research is still required to be performed.
@@
In this paper, we assume that the heavier CP-even Higgs boson $H$ is the observed SM-like Higgs boson whose properties agree with the LHC measurements. We then perform a detailed Monte Carlo (MC) study on the signal process $pp\to Z^{*}\to hA\to 4\tau\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$ (with $l=e,\mu$ and $\tau_{h}$ representing hadronic $\tau$-decays) and relevant background processes, to examine its feasibility at the LHC. We do so after taking into account relevant theoretical and experimental conditions. In order to improve the experimental significance, we will require two Same-Sign (SS) leptons. Our analysis demonstrates that, by using this signature, the relevant parameter space can be either discovered or ruled out by using the current integrated luminosity at the LHC.
@@
The paper is organised as follows. In the next section, we briefly describe the 2HDM and its Yukawa scenarios, then introduce a few Benchmark Points (BPs) for our MC analysis which pass all present theoretical and experimental constraints. In the following section, we perform a detailed collider analysis of these BPs and examine the potential to discover the aforementioned signature within the 2HDM Type-X scenario. Finally, we present some conclusions.
@@ II The 2HDM
where $\phi_{1}$ and $\phi_{2}$ are the two Higgs doublet fields. By hermiticity of such a potential, $\lambda_{1,2,3,4}$ as well as $m_{11,22}^{2}$ are real parameters, while $\lambda_{5,6,7}$ and $m_{12}^{2}$ can be complex, in turn enabling possible Charge and Parity (CP) violation effects in the Higgs sector. Upon enforcing two minimisation conditions of the potential, $m^{2}_{11}$ and $m^{2}_{22}$ can be replaced by $v_{1,2}$, which are the Vacuum Expectation Values (VEV) of the Higgs doublets $\phi_{1,2}$, respectively. Moreover, the coupling $\lambda_{1,2,3,4,5}$ can be substituted by the four physical Higgs masses (i.e., $M_{{h}},M_{{H}},M_{{A}}$ and $M_{{H}^{\pm}}$) and the parameter $\text{sin}(\beta-\alpha)$, where $\alpha$ and $\beta$ are, respectively, the mixing angles between the CP-even states and the angle related to the VEV values, i.e., $\tan\beta=v_{1}/v_{2}$. Thus, independent input parameters can be taken as $M_{{h}},M_{{H}},M_{{A}}$ and $M_{{H}^{\pm}}$, $\lambda_{6}$, $\lambda_{7}$, $\text{sin}(\beta-\alpha)$, $\text{tan}\beta$ and $m^{2}_{12}$.
If both Higgs doublet fields of the general 2HDM couple to all fermions, the ensuing scenario can induce FCNCs in the Yukawa sector at tree level. As intimated, to remedy this, a $Z_{2}$ symmetry is imposed on the Lagrangian such that each fermion type interacts with only one of the Higgs doublets Glashow:1976nt . As a consequence, there are four possible types of 2HDM, namely Type-I, Type-II, Type-X (or lepton-specific) and Type-Y (or flipped). However, such a symmetry is explicitly broken by the quartic couplings $\lambda_{6,7}$ and softly broken by the (squared) mass term $m^{2}_{12}$. In what follows, we shall consider a CP-conserving (i.e., $m^{2}_{12}$ and $\lambda_{5}$ are real) 2HDM Type-X and assume $\lambda_{6}=\lambda_{7}=0$ to forbid explicit breaking of $Z_{2}$ while also taking $m^{2}_{12}$ to be generally small, thereby preventing large FCNCs at tree level, which are incompatible with experiment.
In general, the couplings of the neutral and charged Higgs bosons to fermions can be described by the Yukawa Lagrangian given by Branco:2011iw
@@
[…]
where $\kappa_{f}^{S}$ ($S=h,H$ and $A$) are the Yukawa couplings of the frmion $f$ in the 2HDM, which are illustrated in Tab. 1 for the Type-X under consideration. Here, $V_{ud}$ refers to a Cabibbo-Kobayashi-Maskawa (CKM) matrix element and $P_{L,R}$ denote the left- and right-handed projection operators, respectively. The coupling of the two CP-even states $h$ and $H$ to the gauge bosons $VV$ ($V=W^{\pm},Z$) are proportional to $\sin(\beta-\alpha)$ and $\cos(\beta-\alpha)$, respectively. If we assume that either $h$ or $H$ can be the observed SM-like Higgs boson, the coupling to gauge bosons is obtained for $h$ when $\cos(\beta-\alpha)\rightarrow 0$ and for $H$ when $\sin(\beta-\alpha)\rightarrow 0$. Therefore, each scenario can explain the 125 GeV Higgs signal at the LHC. Following our works Arhrib:2021xmc ; Wang:2021pxc ; Arhrib:2021yqf ; Li:2023btx , we shall focus in the present paper on the scenario where $H$ mimics the observed signal with mass  $\sim\,125$ GeV (as previously intimated).
@@
$h$   $\cos\alpha/\sin\beta$   $\cos\alpha/\sin\beta$   $-\sin\alpha/\cos\beta$
@@
$A$   $\cot\beta$   $-\cot\beta$   $\tan\beta$
Table 1: Yukawa couplings of the fermions $f=u,d$ and $\ell$ to the neutral Higgs bosons $S=h,H$ and $A$ in the 2HDM Type-X.
@@ III Bounds and constraints in the parameter space scans
$M_{{h}}\in[90,125]~{}(\text{GeV}),\quad M_{{H}}=125~{}(\text{GeV}),\quad M_{{A}}\in[90,125]~{}(\text{GeV})$ (4)
$M_{{H}^{\pm}}\in[90,125]~{}(\text{GeV}),\quad\text{sin}(\beta-\alpha)\in[-0.25,0.05],\quad\text{tan}\beta\in[20,200]$
$m_{12}^{2}=M_{{H}^{\pm}}^{2}\frac{\tan\beta}{1+\tan^{2}\beta}~{}(\text{GeV})^{2}.$
@@
  • Oblique EW parameters ($S$, $T$ and $U$Peskin:1990zt ; Peskin:1991sw , which govern the mass splitting between the Higgs states, with the following measured values ParticleDataGroup:2022pth :
@@
To account for potential additional Higgs bosons, exclusion bounds at $95\%$ CL are enforced using the HiggsBounds-5.9.0 program Bechtle:2020pkv , which systematically checks each parameter point against the $95\%$ CL exclusion limits derived from Higgs boson searches conducted by LEP, Tevatron and LHC experiments.
@@
  • BR$(B\to X_{s}\gamma)=(3.32\pm 0.15)\times 10^{-4}$ HFLAV:2022esi ,
  • ${\rm BR}(B^{0}\to\mu^{+}\mu^{-})_{\text{(LHCb)}}$=$\left(1.2^{+0.8}_{-0.7}\right)\times 10^{-10}$ LHCb:2021awg ; LHCb:2021vsc ,
  • ${\rm BR}(B^{0}\to\mu^{+}\mu^{-})_{\text{(CMS)}}$=$\left(0.37^{+0.75}_{-0.67}\right)\times 10^{-10}$ CMS:2022mgd ,
@@
  • ${\rm BR}(B_{s}\to\mu^{+}\mu^{-})_{\text{ (CMS)}}$=$\left(3.83^{+0.38}_{-0.36}\right)\times 10^{-9}$ CMS:2022mgd ,
  • BR$(B\to\tau\nu)=(1.09\pm 0.20)\times 10^{-4}$ HFLAV:2022esi .
Figure 1: Scatter plot of ${\rm{BR}}(h\to\tau^{+}\tau^{-})\times{\rm{BR}}(A\to\tau^{+}\tau^{-})$ as a function of $M_{{h}}$ and $M_{{A}}$ (left panel) and $M_{{A}}-M_{{h}}$ as a function of $\sin(\beta-\alpha)$ and $\tan\beta$ (right panel).
@@
When the constraint on $a_{\mu}$ is considered, there are two types of contributions to $\Delta a_{\mu}$, which are one-loop corrections from $H$, $A$ and $H^{\pm}$ Chun:2016 , as well as two-loop Barr-Zee contributions with heavy fermions Ilisie:2015 . Only in the large $\text{tan}\beta$ regions $\Delta a_{\mu}$ is large enough to explain the newest FNAL measurements. The surviving parameter space is shown in Fig. 2. Compared with Fig. 1 (left), most parameter spaces have been excluded; in fact, only points with $\text{tan}\beta>130$ are available, and the accessible ($M_{h},M_{A}$) regions are rather sparse.
Figure 2: Scatter plot of $\tan\beta$ as a function of $M_{{h}}$ and $M_{{A}}$ when the $a_{\mu}$ effect is considered in additional to the constraints imposed in the previous figure.
To test the allowed parts of the parameter space, nine BPs are given in the Tab. 2, in terms of the 2HDM Type-X parameters entering our production and decay process, including the corresponding $pp\to Z^{*}\to hA\to 4\tau$ cross sections at the LHC with $\sqrt{s}=14$ TeV. Based on this table, we select two scenarios. For BP1-4, where $\tan\beta$ always below 100 and $\sin(\beta-\alpha)$ is close to but not exactly zero. Here, $a_{\mu}$ remains a concern (till new measurements are performed) For BP5-9, the model is in the alignment-limit scenario, where $\sin(\beta-\alpha)=0$ and $\tan\beta$ is very large. In this scenario, $a_{\mu}$ can be explained well (with the present measurements).
$M_{h}$ $M_{A}$ $M_{H}^{\pm}$ $\tan\beta$ $\sin(\beta-\alpha)$ $\sigma_{4\tau,~{}14~{}{\rm TeV}}$ (fb)
BP1 95.10 121.70 95.02 24.19 $-0.076$ 38.76
BP2 120.10 118.50 119.80 41.25 $-0.003$ 27.89
BP3 99.52 102.30 99.53 60.71 $-0.034$ 52.01
BP4 109.10 89.07 109.00 84.43 $-0.025$ 55.23
BP5 105.70 117.40 105.60 105.30 $-0.001$ 35.62
@@
Table 2: 2HDM Type-X input parameters and Leading Order (LO) cross sections (at the parton level with $\sqrt{s}=14$ TeV) are presented for each BP. The unit of all masses is GeV and we fix the SM-like Higgs boson mass as $M_{H}$ = 125 GeV. (Also recall that $m_{12}$ has been taken as a derived quantity, function of $M_{H^{\pm}}$ and $\tan\beta$.)

### v3

@@ Abstract
We analyse light Higgs scalar and pseudoscalar associated hadro-production in the 2-Higgs Doublet Model (2HDM) Type-X (or lepton-specific) within the parameter space allowed by theoretical self-consistency requirements as well as the latest experimental constraints from the Large Hadron Collider (LHC), precision data and $B$ physics. Over the viable regions of such a scenario, the Standard Model (SM)-like Higgs boson discovered at the LHC in 2012 is the heavier CP-even state $H$. Furthermore, in the Type-X scenario, due to large $\text{tan}\beta$, the lighter Higgs scalar $h$ and the pseudoscalar $A$ mainly decay into two $\tau$ leptons. Therefore, we concentrate on analysing the signal process $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}\to\ell\nu_{\ell}\ell\nu_{\ell}\tau_{h}\tau_{h}$ (where $\ell=e,\mu$ whereas $\tau_{h}$ represents the hadronic decay of the $\tau$) and explore the feasibility of conducting such a search at the LHC with a centre-of-mass energy of $\sqrt{s}~=~\text{14 TeV}$ and a luminosity of $L~=~\text{300}~\text{fb}^{-1}$. To suppress the huge SM background, we confine ourselves to consider the fraction of 2HDM events with two same-sign $\tau$ leptons further decaying into same-sign leptons while the other two $\tau$ leptons decay hadronically. We find that a combination of kinematical selection and Machine Learning (ML) analysis will yield significant sensitivity to this process at the end of LHC Run 3.
@@ I Introduction
The LHC has played an indispensable role in improving our understanding of the microcosmic subatomic structure, particularly with the discovery of a 125 GeV scalar particle by the ATLAS and CMS collaborations in 2012 Aad:2012tfa ; Chatrchyan:2012ufa . This was the last particle predicted by the SM, and its properties, as measured by these experiments, are currently consistent with the SM prediction at the $5~\sigma$ level ATLAS:2022vkf ; CMS:2022dwd . Investigating the Higgs sector at the LHC, where the presence of additional (pseudo)scalars is possible, remains an intriguing task, especially in new physics models with additional doublet fields. Therefore, discovering extra Higgs bosons at the LHC is one of the prime targets of Beyond the SM (BSM) signal searches.
@@
One of the simplest extensions of the SM is the 2HDM, which contains two complex Higgs doublets Lee:1973iz ; Deshpande:1977rw ; Branco:2011iw . After Electro-Weak Symmetry Breaking (EWSB), there are three Goldstone bosons, which are ”eaten” by the $W^{\pm}$ and Z{Z} bosons, while the remaining five degrees of freedom become physical Higgs bosons: two neutral CP-even scalars ($h$ and $H$, with $M_{h}\textless M_{H}$), one CP-odd pseudoscalar ($A$), and two charged Higgs states $H^{\pm}$ (with mixed CP status).
@@
In order to be consistent with the stringent experimental constraints from Flavour Changing Neutral Currents (FCNCs) at the tree level, a $Z_{2}$ symmetry is typically introduced into the Yukawa sector so that each type of fermion couples only to one of the doublets of the 2HDM Glashow:1976nt . Based on the $Z_{2}$ charge assignments of the Higgs doublets, we can define four basic scenarios, known as (Yukawa) types. In the lepton-specific scenario (or Type-X), the additional Higgs doublet $\phi_{2}$ couples only to the lepton sector and not to the quark sector, where the light Higgs boson mainly decays into $\tau^{+}\tau^{-}$ over a wide range of parameter space Aoki:2009ha .
@@
Motivated by recent searches for $\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ by the CMS experiment CMS:2015twz ; CMS:2017dmg ; CMS:2019spf ; CMS:2018ams , we will study the process $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ in the 2HDM Type-X. In this scenario, the $h/A\to\tau^{+}\tau^{-}$ decay could be enhanced, resulting in a sizeable $4\tau$ final state. Therefore, it is worthwhile to examine whether this statement is robust enough after taking into account the effects of parton showering, hadronisation, heavy flavour decays and detector resolution.
@@
The results from the muon g2g-2 experiment collected in 2021, conducted at the Fermilab National Accelerator Laboratory (FNAL), when combined with the old ones from the Brookhaven National Laboratory (BNL), gives:
$a_{\mu}^{\rm exp}=116592061(41)\times 10^{-11},$ (1)
where $a_{\mu}=(g-2){\mu}/2$ is the muon anomalous magnetic moment Albahri2021 . The theoretical prediction can be separated into pure QED, EW and hadronic contributions. The pure QED contribution is calculated at five loops whereas two loops are evaluated in the EW interactions. The hadronic contribution, appearing at $\mathcal{O}$(1) GeV due to Hadronic Vacuum Polarisation (HVP) and Ladronic Light-by-Light (HLbL) scattering, provides most of the uncertainty. Using a data-driven dispersive approach, one obtains $a_{\mu}^{\rm SM}=116591810(43)\times 10^{-11}$, which yields a 4.2 $\sigma$ discrepancy between the experimental data and the theoretical prediction from the SM Aoyama2020 :
$\Delta a_{\mu}=a_{\mu}^{\rm exp}-a_{\mu}^{\rm SM}=(251\pm 59)\times 10^{-11}.$ (2)
[…]
@@
However, one can also explain the muon anomalous magnetic moment with new physics models, e.g., the 2HDM Type-X scenario. Since the couplings of the new Higgs bosons to the SM fermions are proportional to $\text{tan}\beta$, the corresponding contributions to $\Delta a_{\mu}$ can be enhanced Jueid2023 ; Iguro:2023tbk . Although a recent measurement has already shown the possibility of restricting the parameter space of the Type-X model in the large $\text{tan}\beta$ region CMS:CMS23007 , further research is still required.
@@
In this paper, we assume that the heavier CP-even Higgs boson $H$ is the observed SM-like Higgs boson whose properties agree with the LHC measurements. We then perform a detailed Monte Carlo (MC) study on the signal process $pp\to Z^{*}\to hA\to 4\tau\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$ (with $l=e,\mu$ and $\tau_{h}$ representing hadronic $\tau$-decays) and relevant background processes, to examine its feasibility at the LHC.
@@
The paper is organised as follows. In the next section, we briefly describe the 2HDM and its Yukawa scenarios, then introduce a few Benchmark Points (BPs) for our MC analysis that pass all present theoretical and experimental constraints. In the following section, we perform a detailed collider analysis of these BPs and examine the potential to discover the aforementioned signature within the 2HDM Type-X scenario. Finally, we present some conclusions.
@@ II The 2HDM
where $\phi_{1}$ and $\phi_{2}$ are the two Higgs doublet fields. By hermiticity of such a potential, $\lambda_{1,2,3,4}$ as well as $m_{11,22}^{2}$ are real parameters, while $\lambda_{5,6,7}$ and $m_{12}^{2}$ can be complex, thereby enabling possible Charge and Parity (CP) violation effects in the Higgs sector. Upon enforcing the two minimisation conditions of the potential, $m^{2}_{11}$ and $m^{2}_{22}$ can be replaced by $v_{1,2}$, which are the Vacuum Expectation Values (VEV) of the Higgs doublets $\phi_{1,2}$, respectively.
Moreover, the couplings $\lambda_{1,2,3,4,5}$ can be substituted with the four physical Higgs masses (i.e., $M_{{{h}}},M_{{{H}}},M_{{{A}}}$ and $M_{{{H}}^{\pm}}$) and the parameter $\text{sin}(\beta-\alpha)$, where $\alpha$ and $\beta$ are, respectively, the mixing angles between the CP-even states and the angle related to the VEV values, i.e., $\text{tan}\beta=v_{1}/v_{2}$. Thus, independent input parameters can be taken as $M_{{{h}}},M_{{{H}}},M_{{{A}}}$ and $M_{{{H}}^{\pm}}$, $\lambda_{6}$, $\lambda_{7}$, $\text{sin}(\beta-\alpha)$, $\text{tan}\beta$ and $m^{2}_{12}$.
If both Higgs doublet fields of the general 2HDM couple to all fermions, the ensuing scenario can induce FCNCs in the Yukawa sector at tree level. As intimated, to remedy this, a $Z_{2}$ symmetry is imposed on the Lagrangian such that each fermion type interacts with only one of the Higgs doublets Glashow:1976nt . As a consequence, there are four possible types of 2HDM, namely Type-I, Type-II, Type-X (or lepton-specific) and Type-Y (or flipped). However, such a symmetry is explicitly broken by the quartic couplings $\lambda_{6,7}$ and softly broken by the (squared) mass term $m^{2}_{12}$.
In what follows, we shall consider a CP-conserving (i.e., $m^{2}_{12}$ and $\lambda_{5}$ are real) 2HDM Type-X and assume $\lambda_{6}=\lambda_{7}=0$ to forbid the explicit breaking of $Z_{2}$, while also taking $m^{2}_{12}$ to be generally small, thereby preventing large FCNCs at tree level, which are incompatible with experimental data.
In general, the couplings of the neutral and charged Higgs bosons to fermions can be described by the Yukawa Lagrangian Branco:2011iw
@@
$\left(\frac{V_{ud}}{\sqrt{2}v}\bar{u}(m_{u}\kappa_{u}^{A}P_{L}+m_{d}\kappa_{d}^{A}P_{R})dH^{+}+\frac{m_{l}\kappa_{l}^{A}}{\sqrt{2}v}\bar{\nu}_{L}l_{R}H^{+}+\text{h.c.}\right),$ (4)
where $\kappa_{f}^{S}$ ($S=h,H$ and $A$) are the Yukawa couplings of the fermion $f$ in the 2HDM, which are illustrated in Tab. 1 for the Type-X under consideration. Here, $V_{ud}$ refers to a Cabibbo-Kobayashi-Maskawa (CKM) matrix element and $P_{L,R}$ denotes the left- and right-handed projection operators, respectively. The couplings of the two CP-even states $h$ and $H$ to the gauge bosons VVVV ($V=W^{\pm},Z$) are respectively proportional to $\sin(\beta-\alpha)$ and $\cos(\beta-\alpha)$. If we assume that either $h$ or $H$ can be the observed SM-like Higgs boson, the coupling to gauge bosons is obtained for $h$ when $\cos(\beta-\alpha)\rightarrow 0$ and for $H$ when $\sin(\beta-\alpha)\rightarrow 0$. Therefore, each scenario can explain the 125 GeV Higgs signal at the LHC. Following our previous works Arhrib:2021xmc ; Wang:2021pxc ; Arhrib:2021yqf ; Li:2023btx , we shall focus in the present paper on the scenario where $H$ mimics the observed signal with mass  $\sim\,125$ GeV (as previously intimated).
@@
$h$   $\cos\alpha/\sin\beta$   $\cos\alpha/\sin\beta$   $\sin\alpha/\cos\beta$
@@
AA   $\cot\beta$   $\cot\beta$   $\text{tan}\beta$
Table 1: Yukawa couplings of the fermions $f=u,d$ and $\ell$ to the neutral Higgs bosons $S=h,H$ and AA in the 2HDM Type-X.
@@ III Bounds and constraints in the parameter space scans
$M_{{{h}}}\in[90,125]~\text{GeV},\quad M_{{{H}}}=125~\text{GeV},\quad M_{{{A}}}\in[90,125]~\text{GeV}$ (5)
$M_{{{H}}^{\pm}}\in[90,125]~\text{GeV},\quad\text{sin}(\beta-\alpha)\in[-0.25,0.05],\quad\text{tan}\beta\in[20,200]$
$m_{12}^{2}=M_{{{H}}^{\pm}}^{2}\frac{\text{tan}\beta}{1+\tan^{2}\beta}~\text{GeV}^{2}.$
@@
  • Oblique EW parameters ($S$, $T$ and UUPeskin:1990zt ; Peskin:1991sw , which govern the mass splitting between the Higgs states, with the following measured values ParticleDataGroup:2022pth :
@@
To account for potential additional Higgs bosons, exclusion bounds are enforced using the HiggsBounds-5.9.0 program Bechtle:2020pkv , which systematically checks each parameter point against the $95\%$ CL exclusion limits derived from Higgs boson searches conducted by LEP, Tevatron and LHC experiments.
@@
In addition, we have cross-validated the above scan results using the latest HiggsTools framework Bahl:2022uwn , which integrates updated exclusion limits and signal-strength likelihoods from LHC 13 TeV data within a unified interface.
@@
[…]
  • ${\rm BR}(B^{0}\to\mu^{+}\mu^{-})_{\text{(LHCb)}}$$\left(1.2^{+0.8}_{-0.7}\right)\times 10^{-10}$ LHCb:2021awg ; LHCb:2021vsc ,
  • ${\rm BR}(B^{0}\to\mu^{+}\mu^{-})_{\text{(CMS)}}$$\left(0.37^{+0.75}_{-0.67}\right)\times 10^{-10}$ CMS:2022mgd ,
@@
  • ${\rm BR}(B_{s}\to\mu^{+}\mu^{-})_{\text{ (CMS)}}$$\left(3.83^{+0.38}_{-0.36}\right)\times 10^{-9}$ CMS:2022mgd ,
@@
Figure 1: Scatter plot of ${\rm{BR}}(h\to\tau^{+}\tau^{-})\times{\rm{BR}}(A\to\tau^{+}\tau^{-})$ as a function of $M_{{{h}}}$ and $M_{{{A}}}$ (left panel) and $M_{{{A}}}-M_{{{h}}}$ as a function of $\text{sin}(\beta-\alpha)$ and $\text{tan}\beta$ (right panel).
[…]
When the constraint on $a_{\mu}$ is considered, there are two types of contributions to $\Delta a_{\mu}$, which are one-loop corrections from $H$, AA and $H^{\pm}$ Chun:2016 as well as two-loop Barr-Zee contributions with heavy fermions Ilisie:2015 . Only in the large $\text{tan}\beta$ region is $\Delta a_{\mu}$ large enough to explain the latest FNAL measurements. The surviving parameter space is shown in Fig. 2. Compared with Fig. 1 (left), most of the parameter space has been excluded: in fact, only points with $\text{tan}\beta>130$ are available and the accessible ($M_{{{h}}},M_{{{A}}}$) regions are rather sparse.
Figure 2: Scatter plot of $\text{tan}\beta$ as a function of $M_{{{h}}}$ and $M_{{{A}}}$ when the $a_{\mu}$ effect is considered in addition to the constraints imposed in the previous figure.
To test the allowed parts of the parameter space, nine BPs are given in Tab. 2, in terms of the 2HDM Type-X parameters entering our production and decay process, including the corresponding $pp\to Z^{*}\to hA\to 4\tau$ cross sections at the LHC with $\sqrt{s}=14$ TeV. Based on this table, we select two scenarios. For BPs 1–4, where $\text{tan}\beta$ is always below 100 and $\sin(\beta-\alpha)$ is close to but not exactly zero, we note that $a_{\mu}$ remains a concern (till new measurements are performed). For BPs 5–9, the model is in the alignment-limit scenario, where $\sin(\beta-\alpha)=0$ and $\text{tan}\beta$ is very large while $a_{\mu}$ can be explained well (with the present measurements).
$M_{{{h}}}$ $M_{{{A}}}$ $M_{{{H}}^{\pm}}$ $\text{tan}\beta$ $\sin(\beta-\alpha)$ $\sigma_{4\tau,~14~{\rm TeV}}$ (fb)
BP1 95.10 121.70 95.02 24.19 0.076-0.076 38.76
BP2 120.10 118.50 119.80 41.25 0.003-0.003 27.89
BP3 99.52 102.30 99.53 60.71 0.034-0.034 52.01
BP4 109.10 89.07 109.00 84.43 0.025-0.025 55.23
BP5 105.70 117.40 105.60 105.30 0.001-0.001 35.62
@@
Table 2: 2HDM Type-X input parameters and Leading Order (LO) cross sections (at the parton level with $\sqrt{s}=14$ TeV) are presented for each BP. The unit of all masses is GeV and we fix the SM-like Higgs boson mass as $M_{{{H}}}$ = 125 GeV. (Also recall that $m_{12}$ has been taken as a derived quantity, function of $M_{{{H}}^{\pm}}$ and $\text{tan}\beta$.)